OpenStudy (lexy148):

can someone help me solve this problem, A copper alloy contains 80% copper by weight . How much copper is in 2.00 kg of this alloy?

OpenStudy (amistre64):

well, we can find 1 percent simply by dividing by 100 80 of those percents is then just multiplying 1% by 80

OpenStudy (amistre64):

if we have n amount of copper n/100 = 1%, therefore: 80n/100 = 80%

OpenStudy (lexy148):

so would the answer be 1%

OpenStudy (amistre64):

no, the answer would be 80% which i demonstrated a reasonable way to approach it.

OpenStudy (amistre64):

a general memorization is to move a decimal point: 80% = 80/100 = .80 but if you cant recall all the memorization that one must do to memorize tricks, then simply applying the math will do you well

OpenStudy (lexy148):

so the answer would be 80%

OpenStudy (lexy148):

fot this problem

OpenStudy (lexy148):

@Hero can you hep me out with this problem

hero (hero):

@lexy148, show your attempt to solve this using the formula I gave you.

hero (hero):

Feel free to use the "Draw" application to post your steps.

OpenStudy (lexy148):

 like this @Hero